The national poultry producers published the following estimates of demand and supply relationships for plucked chicken:




QD = 60.000 – 10.000P



QS = 20.000P




a) Calculate the perfect competition equilibrium price and quantity that will result in the market for plucked chicken (6marks)




b) If the poultry industry was organized as a cartel, calculate the price and quantity that would maximize profits for cartel members.



(Hint: The Total Revenue function for the cartel would be: TR= 6Q – 0.0001Q2, and the Supply function could also be expressed as follows: P= 0.00005Q) (4.5marks)




c) Comparing your answers in parts a) and b), what is the effect on price and output in this market when the cartel is allowed to operate? What is the effect on consumer surplus (do not calculate, only describe/explain)? (4.5 marks)




1
Expert's answer
2022-03-13T18:53:28-0400

QD = 60.000 – 10.000P

QS = 20.000P

a) At Equilibrium, "Q_s=Q_d"

20,000P= 60,000-10,000P

30,000P= 60,000

P= 2

Q*= 20,000(2)= 40,000

b) Cartel

It will operate like Monopoly

"TR= 6Q \u2013 0.0001Q^2"

MR= 6-0.0002Q

C=0.00005

MC= 0.00005

0.00005= 6-0.0002Q

0.0002Q= 6.00005

Q= 30,000

P= 30000(0.00005)= 15

c) The price offered increases while the quantity has reduced. The in turn reduces the consumer surplus.